25 February 2013

The first day at the PEARL Ridge Lab and 0PAL

Today the team visited PEARL for the first time and started setting up the instruments. The temperature at the PEARL Ridge Lab was -36 °C, with only a mild wind. This was 10 °C warmer than the Eureka Weather Station (a fact much appreciated by the team)! Our first drive up to the PEARL Ridge Lab involved two trucks, since the water truck was at the Weather Station. Despite the added room this allowed, another two trips with the truck were needed to bring all the equipment up to the Ridge Lab.

Everyone made good progress preparing their instruments for measurements. While the Sun was technically "up" for about five hours, it was only above the nearby mountains for a couple hours (and barely above, at that). We all hope to start taking measurements tomorrow.

Joseph and Dan began by installing a new chiller for the Bruker. The old one had begun to leak after the many challenges of last year's ACE validation campaign. The installation went smoothly. Next, they plugged in and restarted the Bruker 125HR, which had been winterized in October, 2012. This involved opening the compartments, cleaning around the edges of each panel, and carefully cleaning the scanning mirror rods. A minor issue occurred during the start-up procedures when the beamsplitter was placed into the instrument. While it connected with the instrument, it didn't do so with sufficient pressure to create a secure contact. After consulting with Jim and Paul, the component connecting the beamsplitter compartment to its door was tightened. A solid connection was made, and we continued ensuring all was well with the Bruker. As long as the weather allows for enough sunlight, the Bruker instrument lineshape tests and solar measurements will be performed tomorrow. Debora unpacked her PARIS-IR boxes, finding that everything arrived in good condition. At first, PARIS's computer would not boot. Fortunately, it was fixed by removing the memory card. Debora was also able to align the pick-off mirror with the little bit of sunlight that remained once everything was set up. She also took four background measurements, which looked great. PARIS seems to be ready to measure tomorrow!

Over in the UV-Vis lab, Xiaoyi had great success preparing his three instruments for measurements. PEARL-GBS is measurement-ready. Xiaoyi switched it from direct-moon measurement mode to a combination mode in which it will take zenith-sky/direct-sun/MAX-DOAS measurements during daytime, and perform direct-moon measurements at night. He also reassembled the UT-GBS, installed its new 10 meter fibre bundle, and has begun a series of standard tests. He expects it will be taking measurements soon. Perhaps late tomorrow, if all goes well. SAOZ was brought up to the PEARL roof, and will begin measurements tomorrow. Also at the Ridge Lab, Volodya unpacked the two York University boxes with the tracker, SPS and MAESTRO and tested all components inside. Everything worked fine after setting up the correct COM ports. He also helped other team members with various tasks including fixing a computer and an instrument. This morning at 0PAL, Zen powered down and disassembled the E-AERI, whose MCT detector had stopped working in November. There were many possible reasons for this trouble, with varying degrees of complexity and severity, and several replacement parts were brought up. The simplest case involved the analogue-to-digital (ADC) converter. With the instrument opened up, he easily replaced the ADC module according to the manufacturer's instructions. After reassembling the instrument and powering it up, the E-AERI appeared to function normally. A nominal centre-burst was output, and Zen began taking measurements this afternoon. The first few spectra look good. This quick resolution to a troubling situation was a substantial relief to Zen!

There are a few other issues Zen will work on over the next few days. The hatch isn't closing quite right, and some of the instrument parameters are out of normal ranges. The E-AERI needs some general maintenance, and Zen will take care of this while he is here for the campaign. Otherwise, he gives Emily a hand with the lidar whenever possible.

Also at 0PAL, Emily worked on preparing the CANDAC Raman-Mie-Rayleigh Lidar (CRL) for operation. First, she filled the instrument with distilled water and checked for leaks. Next, she found leaks spouting water everywhere, fixed that, and checked again. After fixing another leak, she found one of the plumbing connectors needed to be replaced with a spare (it too was leaking). At last, all the leaks were found and fixed. Moving forward with the setup, Emily and Chris (working remotely from Halifax) tested and confirmed that Chris has the ability to remotely access and control the laser. Wonderful! They will be able to coordinate and work together again, despite being many thousands of kilometers apart. Lastly, a new challenge emerged. The laser wouldn't flash. Emily has installed a flashlamp, and all had seemed ready to go. It turned out that the CRL is missing a tiny, but vital, component: a wire approximately 5 cm long. Emily is searching for a spare.

ACE ozonesonde flights are scheduled to begin on Thursday, February 28.

It has been a busy day here in Eureka, at the PEARL Ridge Lab, and at 0PAL. The team is relaxing in the lounge, and Dan is writing reports, sorting photos, and preparing to launch the 2013 campaign website. The website will be updated daily with photo albums, so make sure to take a look and share our Canadian Arctic science adventure!

Dan

Back to Daily Reports
Back to Home